OpenStudy (anonymous):

Find the solutions of the function by completing the square. If this method cannot be used explain why. I already did it I need help.

OpenStudy (anonymous):

y = 0.05x^2 -x +1 20y=x^2 - 20x + 20 20y-20=x^2-20x 20y+80= x^2 - 20x + 100 20y+80=(x - 10)^2 20y=(x-10)^2-80 y=(0.05)(x - 10)^2-4

OpenStudy (anonymous):

I completed the square what does it want from me?

OpenStudy (anonymous):

jdoe?

OpenStudy (jdoe0001):

well, finding the solutions means, finding the "roots" or zeros of the quadratic equation

OpenStudy (jdoe0001):

so you'd set y = 0 and solve for "x", to get the 2 roots

OpenStudy (anonymous):

ok...

OpenStudy (jdoe0001):

\(\bf y=(0.05)(x - 10)^2-4 \implies y = \cfrac{5}{100}(x - 10)^2-4\\ \textit{setting y = 0}\\ 0 = \cfrac{5}{100}(x - 10)^2-4 \implies 4= \cfrac{5}{100}(x - 10)^2\\ 80=(x - 10)^2 \implies \pm\sqrt{80} = x-10\\ ---------------------\\ 80 = 2 \times 2 \times 2\times 2\times 5\implies 2^4\times 5 \implies (2^2)^2 \times 5\\ ---------------------\\ \pm\sqrt{80} = x-10 \implies \pm\sqrt{(2^2)^2 \times 5} = x-10 \implies \pm4\sqrt{5}= x-10\)

OpenStudy (anonymous):

holy crap!

OpenStudy (jdoe0001):

and surely you can get it from there

OpenStudy (anonymous):

uhhh

OpenStudy (anonymous):

so it cannot be used?

OpenStudy (anonymous):

Im so confused

OpenStudy (jdoe0001):

well, what would be the values for "x" from there?

OpenStudy (anonymous):

the last line you have

OpenStudy (jdoe0001):

the last line isn't simplified yet

OpenStudy (anonymous):

i don't know :(

OpenStudy (jdoe0001):

well, anyhow, the completion was fine btw... so \({\bf 4\sqrt{5}= x-10 \implies} \begin{cases}+4\sqrt{5}+10=& x\\ -4\sqrt{5}+10=& x\\ \end{cases}\)

OpenStudy (anonymous):

that's all it's asking??? lol...

OpenStudy (jdoe0001):

yeah, it's asking for the "zeros" or "roots", which would be the "solutions" so using your square completion, we did find them

OpenStudy (anonymous):

ohh thank you so much

OpenStudy (jdoe0001):

yw